OpenStudy (anonymous):

One diagonal of a rhombus has endpoints (-6, 9) and (-2, 1). What are the endpoints of the other diagonal? (2, 8) and (-10, 2) (-10, 17) and (2, 7) (-6, 5) and (-2, 4) (2, 4) and (6, 6)

OpenStudy (anonymous):

slope of given diagonal \[m=\frac{ 1-9 }{-2+6 }=\frac{ -8 }{4 }=-2\] slope of other diagonal =-1/m=1/2 check the slope given by the points by the formula one by one \[slope=\frac{ y2-y1 }{x2-x1 } and get the result.\]

OpenStudy (anonymous):

so whats the answer